OpenStudy (anonymous):

Please help will give medal and fan. (: a. 2(x+15) b. x+15 c. x+15/2 d. x+15/x-7

OpenStudy (anonymous):

\[2x ^{2}+16x-210/2(x-7)\]

OpenStudy (anonymous):

simplify ^

OpenStudy (anonymous):

2x^2 + -89x - 735

OpenStudy (anonymous):

\[2(x ^{2}+8x-105)=2(x^2+15x-7x-105) =2{x(x+15)-7(x+15)} =2(x+15)(x-7)

OpenStudy (anonymous):

i think now you can solve.

OpenStudy (anonymous):

Okay, so, on the numerator we have a common factor of two. So we can take that out and simplify the fraction somewhat: \[\frac{ 2x^2+16x-210 }{ 2(x-7) }\] \[\frac{ 2(x^2+8x-105) }{ 2(x-7) }\] \[\frac{ x^2+8x-105 }{ x-7 }\] We can now possibly factorise the numerator (as you may know how to do). I am going to use the ac, b method. So we need two numbers that will multiply to give -105 and add to give +8: These numbers are -7 and 15: x^2 - 7x +15x -105 x(x-7) + 15(x-7) (x+15)(x-7) Now we can add this back onto our fraction and cancel out the (x-7) \[\frac{ (x+15)(x-7) }{ x-7 }\] \[x+15\] So the answer is x+15

OpenStudy (anonymous):

thanks guys(:

OpenStudy (anonymous):

yw